Find where the function fx) - 34 - 16? - 1922 + 6 is increasing and where it is EXAMPLE 1 decreasing -2000 SOLUTION (x) = 12x please box answers in order if you can thank you!
Answer #1

fca) 374_1673_1922+6 fla= 1223 - 48%-3848 12962_42-32) = 121(2-8217+4) sign This depends on the qualch of the three factors o